OpenStudy (anonymous):

5x+4=-6,because x=-2 if then statement

OpenStudy (anonymous):

If 5x+4=-6, then x=-2

Directrix (directrix):

5x+4=-6,because x=-2 The initiating condition that causes 5x +4 to equal -6 is that x = -2. So, in the p implies q form, the conditional would read: If x = -2, then 5x + 4 = -6. @PrincessJerniece